Archive for the ‘Status’ Category

Maintenance affecting London AS on NOC and UKNOC 21/05/2019 – Completed

We will be performing maintenance on the following Aggregation Services AS1, AS2, AS3, AS5, AS7, AS10, AS11 and AS16. Date: 21/05/2019 Time: 22:00 Duration: 30 min Impact: customers on these Aggregation Services will experience a short disruption to service as routers reconnect to the service.


NOC Update to US NOC 07/05/2019 Completed

The latest NOC version 2.50.5 will be deployed to the USNOC. This update includes the following fix: Fix issue related to reverse DNS entries for IP addresses not being wiped when an address is removed from a community. Date: 07/05/2019 Time: 06:15 EDT Duration: 5min Impact: Customers may experience a short disruption to service as…

Read More


AS6 Emergency Maintenance 25/04/2019 – Completed

We have received a software patch from Development to assist with the troubleshooting of the stability of AS6. This patch will be loaded this evening and testing will be performed to provide Development with more information. Date: 25/04/2019 Time: 22:00 Duration: 1 hour Impact: Customers will experience a a few short disruptions to service as…

Read More


Maintenance affecting UK Aggregation Services 10/4/2019 Completed

We will be performing maintenance on the following Aggregation Services in the UK. AS1, AS2, AS3, AS4, AS5, AS8, AS9, AS13, AS16, AS17 Date: 10/4/2019 Time: 22:00 Duration: 30min Impact: customers will experience a short break in connectivity while the service restarts. Disruption will be less than a minute.


New York Maintenance 4/8/2019 Completed

We will be performing network maintenance in New York on the 4th April at 18:00EDT. During this maintenance there will be a short disruption to service. Date: 4/8/2019 Time: 18:00EDT Duration: 30 minutes Impact: we expect a short break in connectivity for no longer than 10 minutes during the maintenance.


AS6 problematic 4/4/2019 – Resolved

08:50 We are aware that some customers on AS6 are experiencing connectivity issues. We are investigating and will provide updates. 09:05 Service restarted due to a stuck process consuming resources. Service has returned to normal.